# Focus: Efficient Keyframe Selection for Long Video Understanding

### v1

@@ Abstract
We propose Focus, Frame-Optimistic Confidence Upper-bound Selection, a training-free, model-agnostic keyframe selection module that selects query-relevant frames under a strict token budget. Focus formulates keyframe selection as a combinatorial pure-exploration (CPE) problem in multi-armed bandits: it treats short temporal clips as arms, and uses empirical means and Bernstein confidence radius to identify informative regions while preserving exploration of uncertain areas. The resulting two-stage exploration-exploitation procedure reduces from a sequential policy with theoretical guarantees, first identifying high-value temporal regions, then selecting top-scoring frames within each region On two long-video question-answering benchmarks, Focus delivers substantial accuracy improvements while processing less than 2% of video frames. For videos longer than 20 minutes, it achieves an 11.9% gain in accuracy on LongVideoBench, demonstrating its effectiveness as a keyframe selection method and providing a simple and general solution for scalable long-video understanding with MLLMs. Code is available at https://github.com/NUS-HPC-AI-Lab/FOCUS.
@@ 1 Introduction
Recent advances in large language models (LLMs) and multimodal LLMs (MLLMs) have significantly improved visual understanding and reasoning. In current frameworks, images are encoded into visual tokens aligned with text and jointly processed by the LLM. Extending this paradigm to videos—especially long, untrimmed ones—introduces a key challenge: the sheer number of frames leads to an overwhelming number of visual tokens, making inference computationally prohibitive.
A common solution is aggressive downsampling (Wang et al., 2022b; Lin et al., 2023; Maaz et al., 2024; Zhang et al., 2025c), but uniformly sampling a handful of frames (e.g., 64 from a one-hour video) often misses critical content (Tang et al., 2025; Zhang et al., 2025b). Increasing the frame rate, on the other hand, causes token explosion (Wang et al., 2024c). This trade-off motivates the need for keyframe selection: choosing a small set of informative frames that preserve semantics while staying within token limits.
Recent methods address this by scoring frame relevance with pre-trained vision-language encoders (e.g., CLIP (Radford et al., 2021) or BLIP (Li et al., 2022)) and then pick the highest-relevance frames (Tang et al., 2025; Zhang et al., 2025b). These text–image matching approaches are typically training-free and plug in easily before the visual encoder in MLLM stacks, retrieving frames with higher relevance other than uniform sampling. Despite their success, current keyframe selection methods still face scalability and efficiency limitations. For a one-hour video at 30 fps (over $10^{5}$ frames), exhaustively scoring all frames entails on the order of $10^{11}$-$10^{12}$ FLOPs with a vision-language encoder like BLIP (Li et al., 2022). This scaling pressure forces existing methods to uniformly sample the video to lower frame rate before the scoring process. This pre-filtering process before keyframe selection undermines the goal of identifying most informative keyframes from all frames (Zhang et al., 2025b; Tang et al., 2025).
@@
The key insight behind Focus is grounded in the observation that natural videos exhibit strong temporal locality: adjacent frames are highly correlated in appearance and motion (Wiegand et al., 2003; Wang et al., 2016; 2022b). This local smoothness naturally extends to frame–query relevance scores. As illustrated in Figure 1, we compute the autocorrelation function (ACF) of relevance scores $r_{t}$ on LongVideoBench and VideoMME. The results show a strong local correlation structure, with a half-life of approximately 5 seconds. This observation implies that exhaustive scoring of all frames is unnecessary. Instead, we can formulate keyframe selection as a bandit problem to adaptively allocate computation: quickly filtering out irrelevant temporal regions, concentrating scoring on promising segments, and ultimately prioritizing the most informative keyframes.
Focus first partitions the video into short temporal clips, each treated as an arm in a multi-armed bandit. The clip selection is then framed as a Combinatorial Pure-Exploration (CPE) problem: the goal is to identify a subset of arms that maximizes expected cumulative relevance under a strict token budget. Each arm maintains an empirical mean relevance and a Bernstein-style confidence radius. Computation is adaptively allocated to clips that are either promising (high mean) or uncertain (large confidence radius), following an optimism-in-the-face-of-uncertainty principle. This iterative process enjoys theoretical convergence guarantees. To leverage parallel computation without sacrificing optimism, we reduce the iterative strategy to a coarse-to-fine schedule: optimistic means guide exploration, while unbiased empirical means inform final arm selection. Within each selected arm, we extract the top-relevance frames to construct the final keyframe set.
We validate the effectiveness of our approach on two video understanding benchmarks, including LongVideoBench (Wu et al., 2024) and Video-MME (Fu et al., 2025). The proposed Focus is tested as an off-the-shelf module on with four popular MLLMs. Focus improves answer accuracy over state-of-the-art keyframe selection baselines across benchmarks while maintaining lower inference cost. The gains are especially pronounced on long-form videos: for videos longer than 20 minutes on LongVideoBench, Focus delivers a 1.9% accuracy improvement while still cutting inference cost.
@@ 2.1 Problem Formulation
Let a video be $V=({\bm{x}}_{1},\ldots,{\bm{x}}_{T})$ and denote the corresponding text query as $q$. Let the frame index set be ${\mathbb{T}}=\{1,\ldots,T\}$. A downstream multimodal LLM $\Phi$ consumes a subset of frames indexed by ${\mathbb{K}}\subseteq{\mathbb{T}}$ with $|{\mathbb{K}}|=k$ and produces an answer $\hat{a}=\Phi\big(q,\{{\bm{x}}_{t}\}_{t\in{\mathbb{K}}}\big)$. Let $R_{\Phi}({\mathbb{K}}\mid V,q)$ denote the task-level utility of the selected frames (e.g., quality of generated answer, relevance to query, or other performance metrics).
@@
However, estimating the contribution of each frame $t$ to the task-level utility is also intractable. We therefore posit that $y_{t}$ is indirectly observable via a vision-language encoder $\psi$ that outputs a relevance score $r_{t}=\psi({\bm{x}}_{t},q;{\bm{\theta}})=y_{t}+\epsilon_{\psi}$, where $\epsilon_{\psi}$ denotes encoder-induced noise. We assume $\epsilon_{\psi}$ follows some distribution that are supported on $[0,1]$ and with zero mean and $\sigma_{\psi}^{2}$ variance. Under this assumption, the relevance score $r_{t}$ is a unbiased estimator of $y_{t}$ which is also commonly used in many works (Tang et al., 2025; Yu et al., 2024) implicitly.
@@ 2.2 Clip-Level Selection as Multi-Armed Bandit
For a video $V=({\bm{x}}_{1},\ldots,{\bm{x}}_{T})$, we partition the timeline into $M$ non-overlapping fixed-length clips $\mathcal{A}=\{A_{a}\}_{a=1}^{M}$, where each clip $A_{a}\subseteq{\mathbb{T}}$ spans frames $s_{a},e_{a}]$ and is treated as a bandit arm. We assume that frame-level utility within the same arm share the same distribution: $y_{t}\sim\mathcal{\nu}_{a}$ for all $t\in[s_{a},e_{a}]$, where $\mathcal{\nu}_{a}$ has mean $\mu_{a}$ and variance $\sigma_{a}^{2}$. We define pulling the arm $a$ as randomly sampling one frame from that clip and observing its query relevance score $r_{t}$ as a reward.
@@ 2.3 Optimistic Confidence Upper-bound Arm Selection
11:     $\tilde{\mu}_{n}(a)\leftarrow\hat{\mu}_{n}(a)-\beta_{a}(n)$
@@
13:     $\tilde{\mu}_{n}(a)\leftarrow\hat{\mu}_{n}(a)+\beta_{a}(n)$
@@
12:Update empirical means $\hat{\mu}_{a}$ for $a\in{\mathbb{A}}_{\text{coarse}}$
@@
Generally, we play a exploration game by pulling an arm $a$ and observing the reward $r_{t}$ at each round $n$. We maintain two core empirical statistics for each arm $a$ during this process: an empirical mean $\hat{\mu}_{a}$ and an empirical Bernstein confidence radius (variance-adaptive) $\beta_{a}$, following the UCV-V style bound (Audibert et al., 2009):
@@
Here $N_{a}(n)$ is the number of pulls for arm $a$ at round $n$ and $n=\sum_{a\in\mathcal{A}}N_{a}(n)$ is the total number of pulls. The confidence radius ensures that the empirical mean is within the confidence radius of the true mean with high probability, i.e.,
$\mathcal{P}\left[\left|\hat{\mu}_{a}-\mu_{a}\right|\leq\beta_{a}(n)\right]\geq 1-\frac{6}{n}.$ (6)
@@
As shown in Algorithm 1, the optimistic confidence upper-bound arm selection starts with an initialization phase where we pull each arm for $q$ times and observe the relevance scores as rewards. We then update the empirical means $\hat{\mu}_{a}$ and compute the confidence radius $\beta_{a}(n)$ for each arm $a$. Note the relevance score $r_{t}$ is an unbiased estimator of $y_{t}$ so we have $\mathbb{E}[\hat{\mu}_{a}]=\mu_{a}$. Then we choose the best $m$ arms using the empirical means $\hat{\mu}_{a}$, i.e., ${\mathbb{A}}_{n}=\mathrm{TopM}(\hat{{\bm{\mu}}},m)$, where $\hat{{\bm{\mu}}}$ is the vector of all arms’ empirical means and $\mathrm{TopM}(\cdot,m)$ returns a set of the $m$ arms with the largest empirical means.
We further refine the arm selection by evaluating the "potential" of each arm. To be specific, for arm $a\in{\mathbb{A}}_{n}$, we compute the lower confidence bound of the empirical mean, i.e., $\mathrm{LCB}_{a}(n)=\hat{\mu}_{a}-\beta_{a}(n)$; for arm $a\notin{\mathbb{A}}_{n}$, we compute the upper confidence bound of the empirical mean, i.e., $\mathrm{UCB}_{a}(n)=\hat{\mu}_{a}+\beta_{a}(n)$. If
@@
Using the optimistic scores $\tilde{\mu}a=\hat{\mu}a+\beta_{a}(n)$, we select the top $\alpha m$ arms, $\mathcal{A}_{\text{coarse}}=\mathrm{TopM}(\tilde{{\bm{\mu}}},,\alpha m)$, and allocate an additional $z$ pulls to each $a\in\mathcal{A}_{\text{coarse}}$ (performed in a single batch). Here, $\alpha$ is a hyperparameter that controls the ratio of the coarse exploration budget to the fine-grained exploration budget. This stage is a batched counterpart of the iterative loop in Algorithm 1: it implements the “optimism in the face of uncertainty” principle by concentrating samples on arms with the largest UCB values, while avoiding per-step scheduling overhead.
@@
After the fine exploitation, we form the final set by selecting the best $m$ arms according to the unbiased empirical means, ${\mathbb{A}}_{\text{fine}}=\mathrm{TopM}(\hat{{\bm{\mu}}},m)$. This choice mirrors $\delta$-PAC identification routines, where optimistic scores guide exploration but the recommendation itself is based on $\hat{\mu}_{a}$ rather than $\tilde{\mu}_{a}$.
@@ 3.1 Experimental Setup
We follow the LMMs-Eval framework Zhang et al. (2024a) and adopt the open-source evaluation protocol from AKS for benchmarks, prompts, and scoring. Our experiments focus on two long-video multiple-choice QA benchmarks: LongVideoBench Wu et al. (2024) and VideoMME Fu et al. (2025). These datasets feature videos lasting up to an hour, where effective keyframe selection becomes crucial for performance. To ensure fair comparison (Tang et al., 2025), we disable subtitles, perform zero-shot evaluation, and keep model parameters frozen—varying only the frame selection strategy (our method versus uniform sampling).
@@
We test both open-source video MLLMs (Qwen2VL (Wang et al., 2024a), LLaVA-OV (Li et al., 2025), LLaVA-Video (Zhang et al., 2025c) and Qwen2-7B (Yang et al., 2024) language model) and the commercial GPT-4o (0513). For frame relevance scoring, we use BLIP ITM (Li et al., 2022) to compute $r_{t}=\psi({\bm{x}}_{t},q;{\bm{\theta}})$, where $r_{t}$ estimates the latent frame-level utility as described in Section 2.1, which is justified as a promising choice by Tang et al. (2025).

### v2

@@ Abstract
We propose Focus, Frame-Optimistic Confidence Upper-bound Selection, a training-free, model-agnostic keyframe selection module that selects query-relevant frames under a strict token budget. Focus formulates keyframe selection as a combinatorial pure-exploration (CPE) problem in multi-armed bandits: it treats short temporal clips as arms, and uses empirical means and Bernstein confidence radius to identify informative regions while preserving exploration of uncertain areas. The resulting two-stage exploration-exploitation procedure reduces from a sequential policy with theoretical guarantees, first identifying high-value temporal regions, then selecting top-scoring frames within each region. On two long-video question-answering benchmarks, Focus delivers substantial accuracy improvements while processing less than 2% of video frames. For videos longer than 20 minutes, it achieves an 11.9% gain in accuracy on LongVideoBench, demonstrating its effectiveness as a keyframe selection method and providing a simple and general solution for scalable long-video understanding with MLLMs. Code is available at https://github.com/NUS-HPC-AI-Lab/FOCUS.
@@ 1 Introduction
Recent advances in large language models (LLMs) and multimodal large language models (MLLMs) have significantly improved visual understanding and reasoning. In current frameworks, images are encoded into visual tokens aligned with text and jointly processed by the LLM. Extending this paradigm to videos—especially long, untrimmed ones—introduces a key challenge: the sheer number of frames leads to an overwhelming number of visual tokens, making inference computationally prohibitive.
A common solution is aggressive downsampling (wang2022internvideo; lin2023video; maaz2023videochatgpt; zhang2024video), but uniformly sampling a handful of frames (e.g., 64 from a one-hour video) often misses critical content (tang2025adaptive; zhang2025q). Increasing the frame rate, on the other hand, causes token explosion (wang2024internvideo2). This trade-off motivates the need for keyframe selection: choosing a small set of informative frames that preserve semantics while staying within token limits.
Recent methods address this by scoring frame relevance with pre-trained vision-language encoders (e.g., CLIP (radford2021learning) or BLIP (li2022blip)) and then pick the highest-relevance frames (tang2025adaptive; zhang2025q). These text-image matching approaches are typically training-free and plug in easily before the visual encoder in MLLM stacks, retrieving frames with higher relevance other than uniform sampling. Despite their success, current keyframe selection methods still face scalability and efficiency limitations. For a one-hour video at 30 fps (over $10^{5}$ frames), exhaustively scoring all frames entails on the order of $10^{11}$-$10^{12}$ FLOPs with a vision-language encoder like BLIP (li2022blip). This scaling pressure forces existing methods to uniformly sample the video to lower frame rate before the scoring process. This pre-filtering process before keyframe selection undermines the goal of identifying most informative keyframes from all frames (zhang2025q; tang2025adaptive).
@@
The key insight behind Focus is grounded in the observation that natural videos exhibit strong temporal locality: adjacent frames are highly correlated in appearance and motion (wiegand2003overview; wang2016temporal; wang2022internvideo). This local smoothness naturally extends to frame-query relevance scores.
Concretely, for each video-query pair we compute a frame-level relevance sequence $\{r_{t}\}$, where $r_{t}$ is the cosine similarity between the visual embedding of frame $t$ and the text embedding of the query produced by BLIP. We then measure temporal dependence via the autocorrelation function (ACF) $\rho(\delta)=\mathrm{corr}(r_{t},r_{t+\delta})$ at lag $\delta$ (in seconds), and aggregate $\rho(\delta)$ across videos. As illustrated in Figure 1, both LongVideoBench and Video-MME exhibit strong short-range correlation: the median ACF remains above $0.5$ for roughly the first $5$ seconds.
This observation implies that exhaustive scoring of all frames is unnecessary. Instead, we can formulate keyframe selection as a bandit problem to adaptively allocate computation: quickly filtering out irrelevant temporal regions, concentrating scoring on promising segments, and ultimately prioritizing the most informative keyframes.
Focus first partitions the video into short temporal clips, each treated as an arm in a multi-armed bandit. The clip selection is then framed as a Combinatorial Pure-Exploration (CPE) problem: the goal is to identify a subset of arms that maximizes expected cumulative relevance under a limited budget. Each arm maintains an empirical mean relevance and a Bernstein-style confidence radius. Computation is adaptively allocated to clips that are either promising (high mean) or uncertain (large confidence radius), following an optimism-in-the-face-of-uncertainty principle. This iterative process enjoys theoretical convergence guarantees. To leverage parallel computation, we reduce the iterative strategy to a coarse-to-fine schedule: optimistic means guide exploration, while unbiased empirical means inform final arm selection. Within each selected arm, we extract the top-relevance frames to construct the final keyframe set.
We validate the effectiveness of our approach on two video understanding benchmarks, including LongVideoBench (wu2024longvideobench) and Video-MME (fu2025video). The proposed Focus is tested as an off-the-shelf module on with four popular MLLMs. Focus improves answer accuracy over state-of-the-art keyframe selection baselines across benchmarks while maintaining lower inference cost. The gains are especially pronounced on long-form videos: for videos longer than 20 minutes on LongVideoBench, Focus delivers a 11.9% accuracy improvement while still cutting inference cost.
@@ 2.1 Problem Formulation
Let a video be $V=({\bm{x}}_{1},\ldots,{\bm{x}}_{T})$ and denote the corresponding text query as $q$. Let the frame index set be ${\mathbb{T}}=\{1,\ldots,T\}$. A downstream MLLM $\Phi$ consumes a subset of frames indexed by ${\mathbb{K}}\subseteq{\mathbb{T}}$ with $|{\mathbb{K}}|=k$ and produces an answer $\hat{a}=\Phi\big(q,\{{\bm{x}}_{t}\}_{t\in{\mathbb{K}}}\big)$. Let $R_{\Phi}({\mathbb{K}}\mid V,q)$ denote the task-level utility of the selected frames (e.g., quality of generated answer, relevance to query, or other performance metrics).
@@
However, estimating the contribution of each frame $t$ to the task-level utility is also intractable. We therefore posit that $y_{t}$ is indirectly observable via a vision-language encoder $\psi$ that outputs a relevance score $r_{t}=\psi({\bm{x}}_{t},q;{\bm{\theta}})=y_{t}+\epsilon_{\psi}$, where $\epsilon_{\psi}$ denotes encoder-induced noise. We assume $\epsilon_{\psi}$ follows some distribution that are supported on $[0,1]$ and with zero mean and $\sigma_{\psi}^{2}$ variance. Under this assumption, the relevance score $r_{t}$ is a unbiased estimator of $y_{t}$ which is also commonly used in many works (tang2025adaptive; yu2024frame) implicitly.
@@ 2.2 Clip-Level Selection as Multi-Armed Bandit
For a video $V=({\bm{x}}_{1},\ldots,{\bm{x}}_{T})$, we partition the timeline into $M$ non-overlapping fixed-length clips $\mathcal{A}=\{A_{a}\}_{a=1}^{M}$, where each clip $A_{a}\subseteq{\mathbb{T}}$ spans frames $s_{a},e_{a}]$ and is treated as a bandit arm. We define pulling arm $a$ as uniformly sampling a frame $t\in A_{a}$ and observing its query relevance score $r_{t}$ as the reward. The unseen frame-level utility of the sampled frame is modeled as $y_{t}\sim\nu_{a}$, where $\nu_{a}$ has mean $\mu_{a}$ and variance $\sigma_{a}^{2}$.
@@ 2.3 Optimistic Confidence Upper-bound Arm Selection
11:     $\tilde{\mu}_{a}(n)\leftarrow\hat{\mu}_{a}(n)-\beta_{a}(n)$
@@
13:     $\tilde{\mu}_{a}(n)\leftarrow\hat{\mu}_{a}(n)+\beta_{a}(n)$
@@
12:Update empirical means $\hat{\mu}_{a}(n)$ for $a\in{\mathbb{A}}_{\text{coarse}}$
@@
Generally, we play a exploration game by pulling an arm $a$ and observing the reward $r_{t}$ at each round $n$. We maintain two core empirical statistics for each arm $a$ during this process: an empirical mean $\hat{\mu}_{a}(n)$ and an empirical Bernstein confidence radius (variance-adaptive) $\beta_{a}(n)$, following the UCV-V style bound (audibert2009exploration):
@@
Here $\hat{\sigma}_{a}^{2}$ is the empirical variance of arm $a$, $N_{a}(n)$ is the number of pulls for arm $a$ at round $n$ and $n=\sum_{a\in\mathcal{A}}N_{a}(n)$ is the total number of pulls. The confidence radius ensures that the empirical mean is within the confidence radius of the true mean with high probability, i.e.,
$\mathcal{P}\left[\left|\hat{\mu}_{a}(n)-\mu_{a}\right|\leq\beta_{a}(n)\right]\geq 1-\frac{6}{n}.$ (6)
@@
As shown in Algorithm 1, the optimistic confidence upper-bound arm selection starts with an initialization phase where we pull each arm for $q$ times and observe the relevance scores as rewards. We then update the empirical means $\hat{\mu}_{a}$ and compute the confidence radius $\beta_{a}(n)$ for each arm $a$. Note the relevance score $r_{t}$ is an unbiased estimator of $y_{t}$ so we have $\mathbb{E}[\hat{\mu}_{a}]=\mu_{a}$. Then we choose the best $m$ arms using the empirical means $\hat{\mu}_{a}(n)$, i.e., ${\mathbb{A}}_{n}=\mathrm{TopM}(\hat{{\bm{\mu}}},m)$, where $\hat{{\bm{\mu}}}$ is the vector of all arms’ empirical means and $\mathrm{TopM}(\cdot,m)$ returns a set of the $m$ arms with the largest empirical means.
We further refine the arm selection by evaluating the "potential" of each arm. To be specific, for arm $a\in{\mathbb{A}}_{n}$, we compute the lower confidence bound of the empirical mean, i.e., $\mathrm{LCB}_{a}(n)=\hat{\mu}_{a}(n)-\beta_{a}(n)$; for arm $a\notin{\mathbb{A}}_{n}$, we compute the upper confidence bound of the empirical mean, i.e., $\mathrm{UCB}_{a}(n)=\hat{\mu}_{a}(n)+\beta_{a}(n)$. If
@@
Using the optimistic scores $\tilde{\mu}_{a}(n)=\hat{\mu}_{a}(n)+\beta_{a}(n)$, we select the top $\alpha m$ arms, $\mathcal{A}_{\text{coarse}}=\mathrm{TopM}(\tilde{{\bm{\mu}}},,\alpha m)$, and allocate an additional $z$ pulls to each $a\in\mathcal{A}_{\text{coarse}}$ (performed in a single batch). Here, $\alpha$ is a hyperparameter that controls the ratio of the coarse exploration budget to the fine-grained exploration budget. This stage is a batched counterpart of the iterative loop in Algorithm 1: it implements the “optimism in the face of uncertainty” principle by concentrating samples on arms with the largest UCB values, while avoiding per-step scheduling overhead.
@@
After the fine exploitation, we form the final set by selecting the best $m$ arms according to the unbiased empirical means, ${\mathbb{A}}_{\text{fine}}=\mathrm{TopM}(\hat{{\bm{\mu}}},m)$. This choice mirrors $\delta$-PAC identification routines, where optimistic scores guide exploration but the recommendation itself is based on the empirical means $\hat{\mu}_{a}(n)$ rather than the optimistic means $\tilde{\mu}_{a}(n)$.
@@ 3.1 Experimental Setup
We follow the LMMs-Eval framework zhang2024lmms and adopt the open-source evaluation protocol from AKS for benchmarks, prompts, and scoring. Our experiments focus on two long-video multiple-choice QA benchmarks: LongVideoBench wu2024longvideobench and VideoMME fu2025video. These datasets feature videos lasting up to an hour, where effective keyframe selection becomes crucial for performance. To ensure fair comparison (tang2025adaptive), we disable subtitles, perform zero-shot evaluation, and keep model parameters frozen—varying only the frame selection strategy (our method versus uniform sampling).
@@
We test both open-source video MLLMs (Qwen2VL (wang2024qwen2), LLaVA-OV (li2024llava), LLaVA-Video (zhang2024video) and Qwen2-7B (yang2024qwen2) language model) and the commercial GPT-4o (0513). For frame relevance scoring, we use BLIP ITM (li2022blip) to compute $r_{t}=\psi({\bm{x}}_{t},q;{\bm{\theta}})$, where $r_{t}$ estimates the latent frame-level utility as described in Section 2.1, which is justified as a promising choice by tang2025adaptive. This also ensure a fair comparison setting as the frame-level utility is estimated using the same model.